10. DUtgonals AC and BD of a trapezium ABCD with AB IIDC
intersect each other at 0. Prove that ar(AODJ = ar(BOCJ.

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
Triangles DAB and CAB are on the o c same base AB and between tho same - parallels AB and DC. :. ar(DAB) = ar(CAB) A B ar(OOA) + ar(OAB) = ar(OAB) + ar(COB) ar(AOD) = ar(BOC).